OpenStudy (anonymous):

A student has secure 40% marks to pass.He gets 40 and fails bu 40 marks.Find the max marks?

OpenStudy (anonymous):

please solve my problem

OpenStudy (anonymous):

he gets 40 marks but fail,right? @khorram

OpenStudy (perl):

your question is difficult to read

OpenStudy (perl):

is this the question A student has secure 40% marks to pass.He gets 40 and fails bu 40 marks.Find the max marks?

OpenStudy (anonymous):

is 2 difficult to read

OpenStudy (anonymous):

yes its right marc

OpenStudy (anonymous):

Hello, "A student has to secure 40% marks to pass. He gets 40 and fails by 40 marks. Find the maximum marks?" * Assumption: When you say fails by, I interpret that as 'got incorrect'. \(40 + 40 = 80\) \(40/80 = 0.5\) \(0.5 \times 100 = 50\%\) \(50\% > 40\%\) So the student has passed the test, and the maximum marks possible is 80.

OpenStudy (anonymous):

but the options are a 150 b 225 c 200 d 300 e 175

OpenStudy (anonymous):

Hmm, did you copy the question word for word?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well I can now officially tag to the train of people who also do not understand your question. Sorry, but I cannot see to understand what your question is asking for.

OpenStudy (anonymous):

its in university sample paper of entry test

OpenStudy (perl):

the phrasing is a bit odd. is a 'mark' a correct answer

OpenStudy (perl):

a mark can be right or wrong answer, (check or x mark )?

OpenStudy (anonymous):

@perl British wise, when one says I got full marks, that means he received full points in American terms.

OpenStudy (anonymous):

So: A student has to secure 40% points to pass. He gets 40 and fails by 40 points. Find the maximum points?

OpenStudy (anonymous):

passing marks is 80 because student got 40 and failed by 40 so basically you have to find 80 is the 40% of which no..? so 80/x * 100 = 40 8000/x=40 x=8000/40 x=200=maximum marks

OpenStudy (anonymous):

I got the ans thnkx

OpenStudy (perl):

he has 40 marks right now, and needs another 40 marks to pass , where passing is defined as 40% score Let x = total number of marks ( 40 + 40 ) / x = .40

OpenStudy (anonymous):

Ah, I see.

OpenStudy (anonymous):

Total Marks is often termed as Maximum Marks..

OpenStudy (perl):

oh , since that is the maximum points you can get on a test

OpenStudy (anonymous):

yes.

OpenStudy (perl):

here is a more challenging problem: A candidate who gets 36% marks in an examination fails by 24 marks but another candidate who gets 43% marks gets 18 more marks than the minimum pass marks. Find the maximum marks and the percentage of pass marks.

OpenStudy (ageta):

40+40=80 .40n=80 n=80/.40=200 The maximum marks is 200

OpenStudy (anonymous):

let max marks=y and pass marks=x 36% of y=x-24 43% of y=x+18 solve for x and y then pass percentage.

OpenStudy (anonymous):

A student has to secure 40% marks to pass. He gets 40 and fails by 40 marks. Find the maximum marks?" * Assumption: When you say fails by, I interpret that as 'got incorrect'. 40+40=80 40/80=0.5 0.5×100=50% 50%>40%

OpenStudy (anonymous):

\[\frac{ 36 y }{ 100 }=x-24,36 y=100x-2400 ...(1)\] \[\frac{ 43 y }{ 100 }=x+18\] 43 y=100x+1800 ...(2) subtract (1) from (2) 43 y-36 y=100 x+1800-100x+2400 7 y=4200 y=4200/7=600 from (1) \[43*600=100 x+1800\] 25800=100 x+1800 25800-1800=100 x 24000=100x x=24000/100=240 pass percentage=?